If u=f(x) and v=g(x) then d(uv)/dx is

Grade:11

4 Answers

grenade
8 years ago
its very simple
first take gx constant and diffrentiade and similarly 
take fs constant and diffrentiate gx
grenade
8 years ago
the answer would be f’(x)g(x)+g’(x)f(x)
 approve if useful above too
ng29
8 years ago
its quite simple , rather its product rule for differentiation
anyway  uv=f(x).g(x)
d(uv)/dx = f(x) g,(x) + g(x) f,(x)
it is answer
 
APPROVE IF USEFUL
Gman Namg
8 years ago
hey are you solving integration by parts
but all the way answer is 
d(uv)/dx = f(x) g,(x) + g(x) f,(x)